1) The vapor pressure of liquid antimony is 400 mm Hg at 1.84×103 K. Assuming that its molar heat of vaporization is constant at 115 kJ/mol, the vapor pressure of liquid Sb is _____ mm Hg at a temperature of 1.81×103 K.

2) The normal boiling point of liquid acetone is 329 K. Assuming that its molar heat of vaporization is constant at 29.0 kJ/mol, the boiling point of CH3COCH3 when the external pressure is 1.21 atm is ______ K.
Answer #1

1)
clasius-clepyron equation

ln(P2/P1) = DH0rxn/R[1/T1 - 1/T2]

P1 = ? mmhg, T1= 1.81*10^3 k

P2 = 400 mmhg,    T2 = 1.84*10^3 k

DHvap = 115 kj/mol R = 8.314 j.k-1.mol-1

ln(400/x) = (115*10^3/8.314)*((1/(1.81*10^3))-(1/(1.84*10^3)))

x = P1 = vapor pressure = 353.14 mmhg

2)

clasius-clepyron equation

ln(P2/P1) = DH0rxn/R[1/T1 - 1/T2]

P1 = 1 atm ( boiling point), T1= 329 k

P2 = 1.21 atm ,    T2 = ? k

DHvap = 29 kj/mol R = 8.314 j.k-1.mol-1

ln(1.21/1) = (29*10^3/8.314)*((1/(329))-(1/(x)))

x = T2 = Boiling point of liquid acetone = 335 k